If the magnitude of the resultant force is to be 9kN directed along the positive x axis, determine the magnitude of force T acting on the eyebolt and its angle delta.
1
Expert's answer
2020-09-22T15:35:35-0400

"T=\sqrt{8^2+9^2-2(8)(9)\cos45}"



"T=6.57KN"



"\theta= \frac {\sin 45}{6.57}"



"\theta= 30.6\degree"